Neon is a monatomic gas, in which inter-molecular forces are extremely weak. 1.00 kg


of neon gas has a density of 0.828 kgm-3 at 100 kPa and 293 K. Find the :


(a) average kinetic energy of each gas molecule


(b) total kinetic energy of the gas.


(c) Root-mean-square speed (crms)


If the neon gas is compressed to double its density, while keeping the temperature the


same:


(d) What are the values for (a)-(c) at the new pressure and volume?

1
Expert's answer
2022-01-04T10:23:10-0500

a)


"KE_m=\dfrac12 mv_\text{rms}^2=6.08·10^{-21}\text{ J}."

b)


"KE=\dfrac32 k_bNT=3.03·10^{-19}\text{ J}."


c)


"v_\text{rms}=\sqrt\dfrac{3RT}{M}=605\text{ m/s}."


d)


"KE_m=\dfrac12 mv_\text{rms}^2=6.08·10^{-21}\text{ J}.\\\space\\
KE_g=\dfrac32 k_bNT=3.03·10^{-19}\text{ J}.\\\space\\
v_\text{rms}=\sqrt\dfrac{3RT}{M}=605\text{ m/s}."